OpenStudy (kc_kennylau):

Find the general solution of the following differential equations: (a) y''+6y'+5y=0 (b) y''+2y'+y=0 (c) y''+4y=0 (d) y''+4y'=0

OpenStudy (shamil98):

I just studied this topic yesterday xD (a) r^2 + 6r + 5 = 0 (r+5)(r+1) = 0 r = - 5 , r =- 1 \[y(x) = c_1e^{-5x} + c_2e^{-x}\]

OpenStudy (kc_kennylau):

You don't need to provide steps, I created those questions

OpenStudy (kc_kennylau):

because I learnt differential equations before but I forgot them

OpenStudy (shamil98):

(b) r^2 + 2r + 1 = 0 (r+1)^2 r = -1 \[y(x) = c_1e^{-x} + c_2xe^{-x}\]

OpenStudy (kc_kennylau):

I just forgot how to change from aux equations to solution

OpenStudy (shamil98):

oh, okay! the formula for the general solution is: \[ y = c_1e^{r_1x} + c_2e^{r_2x}\] if you have two roots if you one root it is: \[y = c_1e^{r_1x} + c_2xe^{r_1x}\] if you have a complex root you use euler's equation: \[y = e^{\alpha x} ( c_1\cos \beta x + c_2 \sin \beta x)\] where \[\alpha = -b/2a , \beta = \frac{ \sqrt{4ac-b^2} }{ 2a }\]

OpenStudy (shamil98):

\[y'' = r^2 , y' = r , y = 1\]

OpenStudy (kc_kennylau):

Please just tell me the answer, I can learn from the answer, if I still don't understand I'll find you thank you :)

OpenStudy (shamil98):

Okay. (c) y'' + 4y = 0 r^2 + 4 = 0 \[r = \pm 2i\] \[\alpha = 0 , \beta = 2\] \[y = c_1 \cos 2x + c_2\sin 2x\]

OpenStudy (shamil98):

(d) y'' + 4y' = 0 r^2 + 4r = 0 r(r+4) = 0 r = -4 ,r = 0 \[y = c_1 e^{-4x} + c_2x\]

OpenStudy (kc_kennylau):

@shamil98 Thank you so much <3

OpenStudy (shamil98):

np :)

OpenStudy (kc_kennylau):

I wanna test what would happen if I regard +-2i as normal roots...

OpenStudy (kc_kennylau):

\[y(x)=\Large c_1e^{2ix}+c_2e^{-2ix}=c_1(\cos2x+i\sin2x)+c_2(\cos2x-i\sin2x)\\\Large(c_1+c_2)\cos2x+(c_1i-c_2i)\sin2x=c_3\cos2x+c_4\sin2x\] Oh that makes sense now thank you :)